Vitamin​ D, whether ingested as a dietary supplement or produced naturally when sunlight falls upon the​ skin, is essential for​ strong, healthy bones. The bone disease rickets was largely eliminated during the​ 1950s, but now there is concern that a generation of children more likely to watch TV or play computer games than spend time outdoors is at increased risk. A recent study of

28002800

children randomly selected found

1919​%

of them deficient in vitamin D. Complete parts a through d.

​a) Construct a​ 98% confidence interval for the true percentage of children who are deficient in vitamin D. Select the correct choice below​ and, if​ necessary, fill in the answer boxes within your choice.

A.The​ 98% confidence interval is

left parenthesis nothing % comma nothing % right parenthesis%,%.

​(Round to one decimal place as​ needed.)

B.

The conditions of a confidence interval are not satisfied.

​b) Does the study show that computer games are a likely cause of​ rickets?

A.

​No, because the interval says nothing about causation.

B.

Yes. Since food manufacturers have increased the amount of vitamin D in their products over the​ years, it is likely the case that children playing too many computer games has caused the increase in rickets.

C.​Yes, because the bone disease was largely​ eliminated, and now we are​ 98% confident that

1919​%

of children have the

deficiency.

D.

​No, because the study suggests that it is more likely the case that a change in​ children's diets over the years is the cause for the increase in rickets.
Answer #1

0 solution : Vitamin Di Warthingoched as a dicalony Grippercentis Rezuired information is given by, D-0800 P=19.1. -19 100 01QEFC-1 -1- 98.1 10.98 -0.00 Whato.collo f. 20.01 212= 200.01) hull= 8. 92.63] From & , we get 98.1. of C1=C649]+ C2-8263)*/0.6 Doels the study show that computer? the computer gernes are likely hk hale to show that Cause of ricketts, - i deres Nou be